Vatic Capital. Cachita Haynes works as a currency speculator for Vatic Capital of Los Angeles. Her latest speculative position is to profit from her expectation that the U.S. dollar will rise significantly against the Japanese yen. The current spot rate is ¥122.00122.00 /$.She must choose between the following 9090 -day options on the Japanese yen:

Option   Strike Price (yen/$)   Premium ($/yen)
Put on yen   124 0.00003/¥
Call on yen   124 0.00046/¥

.

a. Should Cachita buy a put on yen or a call on yen?

b. What is Cachita's breakeven price on the option purchased in part a ?

c. Using your answer from part a , what is Cachita's gross profit and net profit (including premium) if the spot rate at the end of 9090 days is ¥140.00140.00 /$?
Answer #1

Speculate that dollar will appreciate and yen will depreciate

Buy a put on yen

Breakeven price=1/124-0.00003 $/yen or 124.46 yen/$

Gross profit=(1/124-1/140)*12500000=11525$

Net profit=11525-0.00003*12500000=11150$